# DeformTime: Capturing Variable Dependencies with Deformable Attention for Time Series Forecasting

### v2

@@ Appendix B Summary of other forecasting models
  • LightTS [54] uses MLP layers as the building blocks to extract both inter-variable and intra-variable dependencies upon time series with different temporal resolutions.
  • DLinear [53] uses a fully connected layer along the temporal dimension with seasonal-trend decomposition to regress the historical values for future predictions.
  • Crossformer [55] is a transformer-based model that divides the input into patches and proposes a two-stage attention layer to capture inter- and intra-variable dependencies.
  • PatchTST [32] is a transformer-based model that takes segmented patches as input tokens for the model. The prediction of each variable is designed to be independent of one another, i.e. with no inter-variable dependencies established.
  • iTransformer [28] is a transformer-based model that embeds the sequence dimension rather than the input variable dimension. It currently provides the SOTA performance on the long-term forecast tasks. [^15] Long-term forecast tasks ranking, github.com/thuml/Time-Series-Library We note that although iTransformer captures inter-variable dependencies, it achieves that by embedding the input along the sequence dimension. This disrupts the temporal structure within each variable, leading to a sub-optimal performance in our empirical evaluation.
  • TimeMixer [44] is an MLP-based model that predicts the seasonal and trend components at different sampling scales and mixes forecasts to form the final prediction. Note that the authors consider variable mixing optional for their model. We opted to deploy the version that conducts variable mixing for the ILI rate forecasting tasks.
@@
Appendix C Supplementary experiment settings
This section is supplementary to section 4.1 in the main paper. All experiments were conducted (in parallel) using a server with 2 NVIDIA A40 GPUs.
C.1 Input variable re-arrangement based on correlation with the target variable
In the ETT and weather forecasting tasks, we rearrange variables (see NAE in section 3.2) based on their linear correlation with the target variable across all time steps preceding the validation set. In the ILI rate forecasting tasks where training sets cover a period of $9$ years prior to the test period, we only use the last $5$ influenza seasons in the training set to obtain correlations. This is because both user search behaviour and search engine characteristics (e.g. automatic search suggestions) evolve, and this approach to feature selection reinforces a weak recency effect that is ultimately beneficial to prediction accuracy [50, 31].
C.2 DeformTime setup
We use $G=4$ variable neighbourhoods throughout (see section 3.2). This setting ($G=4$) also determines the number of group partitions for offset generation and multi-head attention in T-DAB (section 3.4). The learnable amplitude $\alpha\!\in\!\{3,5,7,9\}$ (see sections 3.3 and 3.4) is tuned collectively for both V-DAB and T-DAB. We keep the kernel size $k$ of the interpolation functions $\phi(\cdot)$ and $\psi(\cdot)$ identical to $\alpha$ throughout the experiments. In all experiments, DeformTime has $2$ encoder layers (see section 3.5). For the ETT and weather tasks, we set $r\!=\!1$ in the first encoder layer and select from $r\!\in\!\{6,12,24\}$ time steps in the second encoder layer (see section 3.4). In addition, we select the segmentation length from $\ell\!\in\!\{6,12,24\}$ time steps (see section 3.3) collectively for both encoder layers. Given the more direct weekly temporal resolution in the reported ILI rates and the fact that web search activity trends do have weekly patterns, in the ILI forecasting task, we set $\ell\!=\!7$ time steps (days), and also set $r\!=\!1$ and $r\!=\!7$ time steps in the first and second encoder layer, respectively. The segmentation stride $s$ is kept the same as $\ell$ throughout the experiments.
C.3 Optimisation and hyperparameter settings
The look-back window $L$ is set to $336$ time steps for the ETTh1, ETTh2, and weather forecasting tasks, based on Nie et al. [32]. For influenza forecasting, we set $L$ to $\{28,28,56,56\}$ days for forecasting horizons $H\!=\!\{7,14,21,28\}$ days ahead, respectively. Only for influenza forecasting, we also set the batch size and $d$ to $64$. For the other tasks, these become learnable parameters (see next section). The number of training epochs is set to $100$ for ETTh1, ETTh2, and weather tasks and $50$ for the influenza forecasting task. Neural networks are optimised with Adam using mean squared error loss on all outputs. Note that the output can be a time series (sequence) of one variable (the target) or more variables (target and exogenous) depending on the forecasting method. As an exception, we use MAE loss for all models tuned on the weather forecasting task. Training stops early when the validation error of the target variable at time $t\!+\!H$ does not decrease further (compared to its lowest value) for $5$ consecutive epochs. We then use the model with the lowest validation loss.
For the ILI rate forecasting tasks, all models are trained with a fixed random seed equal to ‘42’. For ETTh1, ETTh2, and weather tasks, we fix the seed to ‘2021’ based on previous work [32, 53, 44], with the exception of iTransformer where we set the seed to ‘2023’ in accordance with the official configuration.
C.4 Hyperparameter validation
[…]
Appendix D Supplementary results
Table D1: ILI rate forecasting accuracy results in England (ILI-ENG) using web search frequency time series as exogenous variables. $\rho$ and $\epsilon\leavevmode\nobreak\ \%$ denote linear correlation and sMAPE, respectively. The best results are in bold font and the second best are underlined.
@@ Appendix D Supplementary results
D.1 Symmetric Mean Absolute Percentage of Error
To compare results across different tasks that have different units, we use the sMAPE metric. We note the general (and correct) perception that occasionally sMAPE may mislead (by over- or under-estimating error), and hence it can only be used for partial insights in conjunction with a more robust error metric. Hence, our main error metric is MAE. For the ILI rate forecasting tasks, we additionally show linear correlation as an established metric in related literature [10, 20, 50].
$\texttt{sMAPE}\!\left(\hat{\mathbf{y}},\mathbf{y}\right)=\frac{100}{n}\sum_{j=1}^{n}\frac{\left|\hat{y}_{j}-y_{j}\right|}{0.5\left(\left|\hat{y}_{j}\right|+\left|y_{j}\right|\right)}\,.$ (12)
In Tables throughout the manuscript, we denote sMAPE by using $\epsilon\leavevmode\nobreak\ \%$.
Table D2: ILI rate forecasting accuracy results in US Region 2 (ILI-US2) using web search frequency time series as exogenous variables. $\rho$ and $\epsilon\leavevmode\nobreak\ \%$ denote linear correlation and sMAPE, respectively. The best results are in bold font and the second best are underlined.
@@ D.1 Symmetric Mean Absolute Percentage of Error
Table D3: ILI rate forecasting accuracy results in US Region 9 (ILI-US9) using web search frequency time series as exogenous variables. $\rho$ and $\epsilon\leavevmode\nobreak\ \%$ denote linear correlation and sMAPE, respectively. The best results are in bold font and the second best are underlined.
@@
D.2 Detailed ILI forecasting results
@@
Complete results (all test periods, models, and forecasting horizons) in the ILI rate forecasting tasks for England (ILI-ENG), US Region 2 (ILI-US2), and US Region 9 (ILI-US9) are presented in Tables D1D2, and D3, respectively. In addition to MAE and sMAPE, we also show the linear correlation (denoted by $\rho$) between estimates and the target variable throughout each test period.
D.3 Seed robustness for DeformTime
We have examined the robustness of DeformTime across different seeds with hyperparameter tuning by re-running the experiments on the ILI-ENG data set across all forecasting horizons and test sets. The results are enumerated in Table D4. The seed used for the results presented in the main paper (‘42’) provided a performance that did not deviate significantly compared to other seeds. Hence, we conclude that DeformTime is robust to random seed initialisation. We note that compared to the average performance across the explored seeds, seed ‘42’ provides a rather conservative estimate.
@@ D.3 Seed robustness for DeformTime
Table D4: Seed control ($5$ seeds) for DeformTime using the ILI-ENG data set across all forecasting horizons ($H$). $\mu$ and $\sigma$ denote the mean and standard deviation of the $5$ obtained MAEs per test period. The results in the main paper were obtained for seed ‘42’.
@@
D.4 Ablation study for DeformTime – Additional information
D.5 Computational complexity and efficiency of DeformTime – Additional information
$\underbrace{\left(C\!+\!1\right)dL\!+\!d}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\small NAE}}}+2\!\underbrace{\left\{\underbrace{n\!\left[(k^{2}\!+\!1)\ell d\!+\!3\ell d^{2}\!+\!3\ell^{2}d\!+\!n\ell dL\right]}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\small V-DAB}}}+\underbrace{r\!\left[(k\!+\!1)\kappa d\!+\!3\kappa d^{2}\!+\!3\kappa^{2}d\right]}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\small T-DAB}}}+2d\!+\!6d^{2}L\right\}}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\small Encoder}}}+\underbrace{2d^{2}L}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\small Decoder}}}\,.$ (13)
V-DAB: $\mathcal{V}=n\!\left[\left(k^{2}\!+\!1\right)\!\ell d+3\ell d^{2}+3\ell^{2}d+n\ell dL\right]\,\,\text{operations}$ (14)
$L\approx n\ell\implies\mathcal{O}\!\left(d^{2}L+dL^{2}\right)$
T-DAB: $\mathcal{T}=r\!\left[\left(k\!+\!1\right)\kappa d+3\kappa d^{2}+3\kappa^{2}d\right]\,\,\text{operations}$ (15)
$L=\kappa r\implies\mathcal{O}\!\left(d^{2}L\right)$
Encoder: $\mathcal{E}=2\!\left(\mathcal{V}+\mathcal{T}+2d+6d^{2}L\right)\,\,\text{operations}$ (16)
$\implies\mathcal{O}\!\left(d^{2}L+dL^{2}\right)$
Decoder: $\mathcal{G}=2d^{2}L\,\,\text{operations}$ (17)
$\implies\mathcal{O}\!\left(d^{2}L\right)\,.$
Figure D2: GPU memory (VRAM) consumption based on: (A) the look-back window ($L$), and (B) the number of input variables ($C\!+\!1$).
Figure D2 depicts the GPU VRAM memory consumption of DeformTime during training compared to other transformer-based models w.r.t. the length of the look-back window ($L$; part A) and the number of input variables ($C+1$; part B). In these experiments, we set both the batch size and the hidden dimension $d$ to $64$. While assessing the impact of $L$, we set $C\!+\!1$ to $32$, and while assessing the impact of the amount of input variables, we set $L$ to $56$. We then present the average memory consumption across $50$ training epochs.
Overall, DeformTime has a relatively small GPU memory footprint. As the length of the look-back window increases (from $28$ to $140$ time steps), iTransformer (lowest) and DeformTime (second lowest) are showing a stable memory consumption that does not exceed 111\!1 GB. We note that the low memory consumption of iTransformer is expected as it encodes the input over the sequence dimension, i.e. from $\mathbb{R}^{L\times(C+1)}$ to $\mathbb{R}^{d\times(C+1)}$, which guarantees consistent GPU memory usage regardless of the value of $L$. In DeformTime, the input partitioning and adaptive transformations within the V-DAB and T-DAB modules reduce computational complexity, leading to reduced memory consumption, especially when $L$ is large. At the same time, the memory footprints of PatchTST and Crossformer linearly increase with $L$.
As we now increase the number of variables from ($32$ to $512$), other methods display an exponential increase in memory consumption (delayed and considerably lower for iTransformer compared to PatchTST and Crossformer), whereas for DeformTime memory consumption is almost unaffected and stable (e.g. $<\!1\!$ GB for DeformTime, but $>\!24\!$ GB for Crossformer). Hence, we argue that our method can handle a greater amount of exogenous predictors more efficiently.
D.6 A brief note about the performance and evaluation of DLinear
@@
Although DLinear achieved competitive results in prior work [53], the performance drop in our paper originates from the fact that we evaluate models based on their estimate for $y_{t+H}$, i.e. the target variable’s value at time step $t\!+\!H$ (see a detailed description of the forecasting task in section 2). In fact, for most tasks this is the target forecast (represented by the corresponding forecasting horizon $H$) and that is what forecasting accuracy should be measured on. Contrary to common sense, DLinear was evaluated on the entire time series (entire sequence of predictions), i.e. from $y_{t+1}$ to $y_{t+H}$. Oddly, this evaluation was conducted in a uniform way, i.e. the error for each time step incurred the same penalty. That may be relevant for some tasks, but it is not relevant to tasks where forecasting $H$ time steps ahead really necessitates obtaining an accurate forecast $H$ time steps ahead (and that should be the case for most, if not all, forecasting tasks). DLinear was obviously more accurate in early time steps (lower degree of difficulty), but very inaccurate in later ones (greater degree of difficulty). Hence, in our experiments DLinear displayed the worst forecasting performance.
D.7 Should MTS forecasting models mix variables in the embedding space?

### v3

@@ Appendix C Summary of other forecasting models
  • LightTS (Zhang et al., 2022) uses MLP layers as the building blocks to extract both inter-variable and intra-variable dependencies upon time series with different temporal resolutions.
  • DLinear (Zeng et al., 2023) uses a fully connected layer along the temporal dimension with seasonal-trend decomposition to regress the historical values for future predictions.
  • Crossformer (Zhang & Yan, 2023) is a transformer-based model that divides the input into patches and proposes a two-stage attention layer to capture inter- and intra-variable dependencies.
  • PatchTST (Nie et al., 2023) is a transformer-based model that takes segmented patches as input tokens for the model. The prediction of each variable is designed to be independent of one another, i.e. with no inter-variable dependencies established.
  • iTransformer (Liu et al., 2024) is a transformer-based model that embeds the sequence dimension rather than the input variable dimension. We note that although iTransformer captures inter-variable dependencies, it achieves that by embedding the input along the sequence dimension. This disrupts the temporal structure within each variable, leading to a sub-optimal performance in our empirical evaluation.
  • TimeMixer (Wang et al., 2024a) is an MLP-based model that predicts the seasonal and trend components at different sampling scales and mixes forecasts to form the final prediction. Note that the authors consider variable mixing optional for their model. We opted to deploy the version that conducts variable mixing for the ILI rate forecasting tasks.
  • TimeXer (Wang et al., 2024b) is a transformer-based model that captures exogenous information with cross-attention. It currently claims to offer SOTA accuracy on longer-term forecasting tasks. [^17] Long-term forecast tasks ranking, github.com/thuml/Time-Series-Library Similarly to iTransformer, TimeXer embeds the exogenous variables along the sequence dimension which disrupts the temporal structure.
  • ModernTCN (Luo & Wang, 2024b) is a CNN-based model that convolutes over both temporal and variable dimensions respectively with large receptive fields to capture inter- and intra-variable dependencies.
  • CycleNet (Lin et al., 2024) is an MLP-based model that disentangles periodic patterns to capture temporal dependencies. Similarly to PatchTST, CycleNet predicts each variable independently. Note that the authors propose two versions of CycleNet. We report results for the version that has $1$ linear layer, which has higher accuracy for overlapping tasks (ETT and weather).
@@
Table S1: ILI rate forecasting results in England (ILI-ENG) using web search frequency time series as exogenous variables. $\rho$ and $\epsilon\leavevmode\nobreak\ \%$ denote linear correlation and sMAPE, respectively. The best results are in bold font and the second best are underlined.
@@
Appendix D Supplementary experiment settings
This section is supplementary to section 4.1 in the main paper. All experiments were conducted using a Linux server with 2 NVIDIA A40 GPUs (and 2 AMD EPYC 7443 CPUs), except for the ablation experiments which were conducted using a Linux server with 3 NVIDIA L40S GPUs (and 2 AMD EPYC 9354 CPUs).
D.1 Input variable re-arrangement based on correlation with the target variable
In the ETT and weather forecasting tasks, we rearrange variables (see NAE in section 3.2) based on their linear correlation with the target variable across all time steps preceding the validation set. In the ILI rate forecasting tasks where training sets cover a period of $9$ years prior to the test period, we only use the last $5$ influenza seasons in the training set to obtain correlations. This is because both user search behaviour and search engine characteristics (e.g. automatic search suggestions) evolve, and this approach to feature selection reinforces a weak recency effect that is ultimately beneficial to prediction accuracy (Yang et al., 2015; Morris et al., 2023).
D.2 Hyperparameters specific to the ILI forecasting task
[…]
D.3 Random seed initialisation
For the ILI rate forecasting tasks, we use a fixed seed equal to ‘42’ during training. For the ETTh1, ETTh2, and weather tasks, we set this to ‘2021’ based on previous work (Nie et al., 2023; Zeng et al., 2023; Wang et al., 2024a), with the exception of iTransformer (‘2023’), and CycleNet, where results are averaged across $5$ seeds (‘2024’, ‘2025’, ‘2026’, ‘2027’, ‘2028’) in accordance with their official configuration.
Table S2: ILI rate forecasting results in US Region 2 (ILI-US2) using web search frequency time series as exogenous variables. $\rho$ and $\epsilon\leavevmode\nobreak\ \%$ denote linear correlation and sMAPE, respectively. The best results are in bold font and the second best are underlined.
@@ D.3 Random seed initialisation
D.4 Data normalisation
In all forecasting tasks, we standardise all variables (zero mean, unit standard deviation), each time based on the training data. The prediction output is de-normalised back to its original scale prior to being compared with the (de-normalised) ground truth. In addition to that, for the ETT and weather data sets, we also standardise each variable within the input’s look-back window, in accordance with previous work (Liu et al., 2022b; Zhou et al., 2022a; Nie et al., 2023).
Appendix E Supplementary results
E.1 Symmetric Mean Absolute Percentage Error
To compare results across different tasks that have different units, we use the sMAPE metric. We note the general (and correct) perception that occasionally sMAPE may mislead (by over- or under-estimating error), and hence it can only be used for partial insights in conjunction with a more robust error metric. Hence, our main error metric is MAE. For the ILI rate forecasting tasks, we additionally show linear correlation as an established metric in related literature (Ginsberg et al., 2009; Lampos et al., 2015; Yang et al., 2015).
$\texttt{sMAPE}\!\left(\hat{\mathbf{y}},\mathbf{y}\right)=\frac{100}{n}\sum_{j=1}^{n}\frac{\left|\hat{y}_{j}-y_{j}\right|}{0.5\left(\left|\hat{y}_{j}\right|+\left|y_{j}\right|\right)}\,.$ (S1)
In Tables throughout the manuscript, we denote sMAPE using $\epsilon\leavevmode\nobreak\ \%$.
Table S3: ILI rate forecasting results in US Region 9 (ILI-US9) using web search frequency time series as exogenous variables. $\rho$ and $\epsilon\leavevmode\nobreak\ \%$ denote linear correlation and sMAPE, respectively. The best results are in bold font and the second best are underlined.
@@ E.1 Symmetric Mean Absolute Percentage Error
E.2 Ablation study for DeformTime – Additional information
E.3 Detailed ILI forecasting results
@@
Complete results (all test periods, models, and forecasting horizons) in the ILI rate forecasting tasks for England (ILI-ENG), US Region 2 (ILI-US2), and US Region 9 (ILI-US9) are presented in Tables S1S2, and S3, respectively. In addition to MAE and sMAPE, we also show the linear correlation (denoted by $\rho$) between estimates and the target variable throughout each test period.
E.4 The effect of variable grouping in the NAE module
Table S4: Ablation study on different groups $G$ for DeformTime using the ILI-ENG and ILI-US9 data sets. Results are averaged across all $4$ seasons.
$G\!=\!2$ $G\!=\!4$ $G\!=\!8$ $G\!=\!16$
$H$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$
ILI-ENG $7$ 1.7733 34.79 1.6417 28.61 1.7325 23.52 1.7085 24.73
$14$ 2.3754 32.92 2.2308 33.98 2.1954 31.52 2.3009 29.50
$21$ 2.7963 35.23 2.6500 32.70 2.7759 43.23 2.8562 34.70
$28$ 3.1379 44.84 2.7228 40.44 2.9543 42.16 3.0996 41.44
$G\!=\!2$ $G\!=\!4$ $G\!=\!8$ $G\!=\!16$
$H$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$
ILI-US9 $7$ 0.2665 12.80 0.2622 12.26 0.2650 11.85 0.2675 12.84
$14$ 0.3253 14.82 0.3084 13.80 0.2926 13.39 0.2995 13.71
$21$ 0.3475 16.27 0.3179 14.24 0.3273 15.35 0.3534 16.12
$28$ 0.3769 16.89 0.3532 15.74 0.3550 15.96 0.3679 16.73
We further explored the effect of the grouping parameter $G$ in DeformTime, which determines both the number of groups in NAE and the number of heads in the multi-head attention of T-DAB. Experiments are conducted using the ILI forecasting task for England (ILI-ENG) and US Region 9 (ILI-US9). Hyperparameters are re-tuned for each value of $G\!\in\!\{2,4,8,16\}$.
Results are enumerated in Table S4. It can be argued that more groups tend to benefit performance for shorter forecasting horizons, whereas a more moderate level of grouping is required for the optimal performance as the forecasting horizon increases. Notably, extensive $(G=16)$ or overly small $(G=2)$ groupings result in performance degradation. This observation could be further interpreted in conjunction with the results presented in Table 3, where we see that DeformTime performs better in shorter-term forecasting tasks when it uses more variables ($C$). Presumably, the model benefits from an adaptive $G$, where $G$ is proportional to the number of input variables $C$. Nevertheless, to reduce the complexity of hyperparameter optimisation in our experiments we have chosen to use a constant value, setting $G\!=\!4$.
We note that the observed influence of grouping may pose a potential limitation of DeformTime, as it affects how dependencies are captured within the latent variables. Specifically, smaller grouping numbers allow the dependencies to be captured among variables holding less linear correlations and vice versa. However, there likely exists a trade-off between allowing more variables to be able to be captured by V-DAB and the model being overfitted as we have empirically shown in Tables 3 and S4, where both removing the NAE and using an overly small number of groups lead to an inferior performance.
E.5 Computational complexity and efficiency of DeformTime – Additional information
$\begin{split}&\underbrace{\left(C\!+\!1\right)dL\!+\!d}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\small NAE}}}+\underbrace{2d^{2}L}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\small Decoder}}}+\\ &2\!\underbrace{\left\{\underbrace{n\!\left[(k^{2}\!+\!1)\ell d\!+\!3\ell d^{2}\!+\!3\ell^{2}d\!+\!n\ell dL\right]}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\small V-DAB}}}+\underbrace{r\!\left[(k\!+\!1)\kappa d\!+\!3\kappa d^{2}\!+\!3\kappa^{2}d\right]}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\small T-DAB}}}+2d\!+\!6d^{2}L\right\}}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\small Encoder}}}\,.\end{split}$ (S2)
V-DAB: $\mathcal{V}=n\!\left[\left(k^{2}\!+\!1\right)\!\ell d+3\ell d^{2}+3\ell^{2}d+n\ell dL\right]\,\,\text{operations}$ (S3)
$L\approx n\ell\implies\mathcal{O}\!\left(d^{2}L+dL^{2}\right)$
T-DAB: $\mathcal{T}=r\!\left[\left(k\!+\!1\right)\kappa d+3\kappa d^{2}+3\kappa^{2}d\right]\,\,\text{operations}$ (S4)
$L=\kappa r\implies\mathcal{O}\!\left(d^{2}L\right)$
Encoder: $\mathcal{E}=2\!\left(\mathcal{V}+\mathcal{T}+2d+6d^{2}L\right)\,\,\text{operations}$ (S5)
$\implies\mathcal{O}\!\left(d^{2}L+dL^{2}\right)$
Decoder: $\mathcal{G}=2d^{2}L\,\,\text{operations}$ (S6)
$\implies\mathcal{O}\!\left(d^{2}L\right)\,.$
@@ E.5 Computational complexity and efficiency of DeformTime – Additional information
Table S5: Seed control ($5$ seeds) for DeformTime using the ILI-ENG data set across all forecasting horizons ($H$). $\mu$ and $\sigma$ denote the mean and standard deviation of the $5$ obtained MAEs per test period. The results in the main paper were obtained for seed ‘42’.
@@
E.6 Seed robustness for DeformTime
We have examined the robustness of DeformTime across different seeds with hyperparameter tuning by re-running the experiments on the ILI-ENG data set across all forecasting horizons and test sets. The results are enumerated in Table S5. The seed used for the results presented in the main paper (‘42’) provided a performance that did not deviate significantly compared to other seeds. Hence, we conclude that DeformTime is robust to random seed initialisation. We note that compared to the average performance across the explored seeds, seed ‘42’ provides a rather conservative estimate.
E.7 A brief note about the performance and evaluation of DLinear
@@
Although DLinear achieved competitive results in prior work (Zeng et al., 2023), the performance drop in our paper originates from the fact that we evaluate models based on their estimate for $y_{t+H}$, i.e. the target variable’s value at time step $t\!+\!H$ (see a detailed description of the forecasting task in section 2). In fact, for most tasks this is the target forecast (represented by the corresponding forecasting horizon $H$) and that is what forecasting accuracy should be measured on. Contrary to common sense, DLinear (as well as other works, e.g. Wang et al. (2024b) or Liu et al. (2024)) was evaluated on the entire time series (entire sequence of predictions), i.e. from $y_{t+1}$ to $y_{t+H}$. Oddly, this evaluation was conducted in a uniform way, i.e. the error for each time step incurred the same penalty. That may be relevant for some tasks, but it is not relevant to tasks where forecasting $H$ time steps ahead really necessitates obtaining an accurate forecast $H$ time steps ahead (and that should be the case for most, if not all, forecasting tasks). The reason behind this is that forecasting errors tend to increase as the forecasting horizon extends and the forecasting task becomes harder (Hyndman & Koehler, 2006). Models may perform differently at short versus long horizons. Averaging the error across all the outputs may favour forecasters that are very accurate early on, but very inaccurate closer to and at the target forecasting horizon. DLinear was obviously more accurate in early time steps (lower degree of difficulty), but very inaccurate in later ones (greater degree of difficulty). Hence, in our experiments, DLinear displayed the worst forecasting performance. The following section (E.8) offers additional insights.
Table S6: Forecasting accuracy results (MAE) using the entire output sequence (as opposed to using the last prediction only at the target forecasting horizon) across all tasks and methods. $H$ denotes the forecasting horizon. For the ILI tasks, errors are averaged across all $4$ test seasons for each region.
Task $H$ DeformTime ModernTCN CycleNet TimeXer PatchTST iTransformer TimeMixer Crossformer LightTS DLinear
ETTh1 96 0.1840 0.1774 0.1746 0.1807 0.1757 0.1825 0.1834 0.1989 0.2089 0.2045
192 0.1988 0.2013 0.1990 0.2043 0.1999 0.2051 0.2049 0.2149 0.2267 0.2834
336 0.2024 0.2123 0.2189 0.2226 0.2248 0.2322 0.2310 0.2577 0.2487 0.3877
720 0.2390 0.2288 0.2367 0.2292 0.2500 0.2449 0.2463 0.3000 0.4450 0.5007
ETTh2 96 0.2924 0.2790 0.2849 0.2797 0.2765 0.2991 0.2963 0.3254 0.3125 0.2840
192 0.3010 0.3183 0.3223 0.3331 0.3174 0.3463 0.3428 0.3534 0.3496 0.3313
336 0.3172 0.3350 0.3536 0.3769 0.3417 0.3857 0.3758 0.3891 0.3939 0.4056
720 0.3577 0.4259 0.3976 0.4069 0.3887 0.3945 0.3940 0.4248 0.4339 0.5391
Weather 96 0.0226 0.0235 0.0228 0.0270 0.0204 0.0218 0.0272 0.0247 0.0263 0.0206
192 0.0237 0.0265 0.0257 0.0294 0.0239 0.0253 0.0309 0.0271 0.0292 0.0239
336 0.0258 0.0296 0.0279 0.0312 0.0259 0.0269 0.0325 0.0311 0.0294 0.0261
720 0.0308 0.0365 0.0333 0.0355 0.0316 0.0312 0.0369 0.0343 0.0339 0.0307
ILI-ENG 7 1.2802 1.7351 2.0997 1.7086 1.9402 1.7746 1.7745 1.3243 1.9713 2.1892
14 1.6521 2.2656 2.6170 1.9703 2.4857 2.3660 2.2577 1.9326 2.3676 2.7452
21 2.3246 2.3331 3.2527 2.8852 2.8458 3.0904 2.9285 2.8733 2.7742 3.2983
28 2.7829 2.5948 3.6091 2.7894 3.2161 3.4784 3.2425 3.0740 3.2559 3.6805
ILI-US2 7 0.4142 0.4523 0.6974 0.4931 0.6400 0.5455 0.5156 0.4444 0.4323 0.7214
14 0.4544 0.5501 0.7601 0.5402 0.7302 0.6357 0.5496 0.4695 0.4723 0.7899
21 0.5374 0.5576 0.8130 0.6190 0.7489 0.6554 0.5966 0.5720 0.6432 0.8642
28 0.5554 0.5892 0.8729 0.6808 0.8450 0.7370 0.6688 0.5842 0.6886 0.8920
ILI-US9 7 0.2614 0.2945 0.4126 0.3435 0.3882 0.3725 0.3322 0.3222 0.3078 0.4348
14 0.3040 0.3398 0.4584 0.3715 0.4334 0.4049 0.3590 0.3254 0.4322 0.4780
21 0.3187 0.3558 0.5224 0.4430 0.4664 0.4504 0.3714 0.3285 0.4385 0.5315
28 0.3913 0.3786 0.5727 0.4774 0.5002 0.5535 0.4236 0.4018 0.5665 0.5622
E.8 Evaluating forecasts across the entire output time series
For a more comprehensive assessment of model performance, we provide forecasting accuracy results where we consider the entire output sequence (the error is averaged across the entire output sequence). Results are enumerated in Table S6. DeformTime outperforms baselines on $17$ out of $24$ tasks, with an averaged MAE reduction by $1.7\%$. When compared to ModernTCN, the best-performing baseline, DeformTime reduces the MAE by $7.2\%$ on average across all tasks.
For the ETT and weather benchmark datasets, the best-performing baseline is PatchTST; DeformTime reduces PatchTST’s MAE by $1.5\%$ on average. If we compare that to obtaining MAE based on the forecasts at the target forecasting horizon only (Table 1), we can see a major discrepancy: DeformTime reduces PatchTST MAE by $9.8\%$ (on the ETT and weather forecast tasks; reduction is greater for the ILI forecasting task) when forecasting accuracy is evaluated based on the actual target forecast. Going back to using the entire output time series, for the ILI tasks, DeformTime reduces the MAE by $3.4\%$ on average. For these tasks, Crossformer achieves the best MAE performance amongst baselines and DeformTime reduces that by $8.2\%$ on average.
Notably, while PatchTST outperforms other baselines when evaluated over the entire sequence in the ETT and weather data sets, CycleNet was the best-performing baseline in our main assessment (Table 1). Hence, by this observation alone, it is evident that while CycleNet is producing more accurate forecasts at the target horizon compared to PatchTST, it yields an inferior average MAE when considering the entire series of forecast outputs from time step $t\!+\!1$ to time step $t\!+\!H$. A similar ranking difference of baseline performance is also observed in the ILI tasks, where ModernTCN was the most competitive baseline when evaluating the results over the target horizon (DeformTime reduces ModernTCN MAE by $11.2\%$ on the ILI tasks as enumerated in Table 1). The overall changes in the ranking of baselines consolidate our argument in section E.7 that convoluting the error at the target forecasting horizon with errors in time steps prior to that can produce a distorted picture of the actual forecasting capacity of a model.
E.9 Should forecasting models mix variables in the embedding space?